# Splitting and gluing in sine-dilaton gravity: matter correlators and the wormhole Hilbert space

### v1

@@ 2.2 EOW brane quantization
$K_{AdS}^{R}=m_{R},\quad\partial_{n_{\mathrm{AdS}}^{R}}\Phi_{\mathrm{AdS}}^{R}=\bar{\mu}+im_{R}e^{-i\Phi/2}.$ (45)
@@
$\mathds{1}=\int_{0}^{\pi}d\theta\leavevmode\nobreak\ \rho(\theta)|\theta\rangle\langle\theta|=\sum_{n=0}^{\infty}|n\rangle\langle n|,\quad\langle n|\theta\rangle=\frac{H_{n}(\cos(\theta)|q^{2})}{(q^{2};q^{2})_{n}^{1/2}}.$ (46)
The only unknown quantity is the wavefunction $\langle n|m_{R},\bar{\mu}\rangle$. As explained in [20], this wavefunction can be determined by expressing the effective mass of the brane in AdS space in two different ways.777 The first way is to directly match the parameter of EOW brane here to the JT gravity brane mass $m_{\mathrm{AdS}}$: $m_{\mathrm{AdS}}=\partial_{n_{\mathrm{AdS}}}\Phi_{\mathrm{AdS}}-\Phi_{\mathrm{AdS}}K_{\mathrm{AdS}}$. The second way is to use the Hamiltonian constraint to relate $m_{\mathrm{AdS}}$ to the JT momentum and extrinsic curvature. See [20] for details.
@@
where an operator ordering has been chosen. We can treat the above equation as an Schrodinger eigenvalue equation
@@
$H_{m_{R},\bar{\mu}}\rightarrow g^{-1}(L)H_{m_{R},\bar{\mu}}\leavevmode\nobreak\ g(L),\quad g(L)=\prod_{k=0}^{\infty}\left(1+2im_{R}e^{-(L+k\hbar)}-e^{-2(L+k\hbar)}\right)^{-1/2},$ (54)
@@
$2\cos(\theta){{}_{\Delta_{R},\bar{\alpha}}\langle}\theta|n\rangle=_{\Delta_{R},\bar{\alpha}}{\langle}\theta|n+1\rangle+(q^{2\Delta}e^{i\bar{\alpha}}+q^{2\Delta}e^{-i\bar{\alpha}})q^{n}{{}_{\Delta_{R},\bar{\alpha}}\langle}|n\rangle+(1-q^{4\Delta}q^{n-1})(1-q^{n}){{}_{\Delta_{R},\bar{\alpha}}\langle}\theta|n-1\rangle.$ (60)
This is exactly the recursion relation for the Al Salam-Chihara polynomial $Q_{n}(\cos(\theta)|q^{2\Delta}e^{i\bar{\alpha}},q^{2\Delta}e^{-i\bar{\alpha}};q^{2})$. Taking into account the fact that the left and right eigenfunctions differ for a non-Hermitian Hamiltonian, the wavefunctions can be normalized as [20, 38]
@@
$\mathds{1}=\int_{0}^{\pi}d\theta\leavevmode\nobreak\ \rho_{\Delta_{R},\bar{\alpha}}(\theta)\,|\theta\rangle_{\Delta_{R},\bar{\alpha}}\,{{}_{\Delta_{R},\bar{\alpha}}}\langle\theta|=\sum_{n=0}^{\infty}|n\rangle\langle n|.$ (65)
@@ 2.3 Gluing condition and the two-point function
$=\int_{0}^{\pi}d\bar{\alpha}\leavevmode\nobreak\ \rho(\bar{\alpha})\langle n=0|e^{-(\beta-\tau)\hat{H}_{\mathrm{SD}}}e^{-\Delta_{L}\hat{L}}|\bar{\alpha}\rangle_{\mathrm{FC}}\langle\bar{\alpha}|e^{-\Delta_{R}\hat{L}}e^{-\tau\hat{H}_{\mathrm{SD}}}|n=0\rangle_{\mathrm{FC}},$ (70)
$=\int_{0}^{\pi}d\bar{\alpha}\leavevmode\nobreak\ \rho(\bar{\alpha})\Psi_{L}(\beta-\tau,\bar{\alpha},\Delta_{L})\Psi_{R}(\tau,\bar{\alpha},\Delta_{R}),$
$=\int_{0}^{\pi}d\bar{\alpha}\leavevmode\nobreak\ \rho(\bar{\alpha})Z_{L}^{\star}(\beta-\tau,\bar{\mu},m_{L})Z_{R}(\tau,\bar{\mu},m_{R}),$ (71)
@@
$\langle\bar{\alpha}|e^{-\Delta_{R}\hat{L}}e^{-\tau\hat{H}_{\mathrm{SD}}}|n=0\rangle=\langle\bar{\alpha}|q^{\Delta_{R}\hat{n}}e^{-\tau\hat{H}_{\mathrm{SD}}}|n=0\rangle=\langle m_{R},\bar{\mu}|e^{-\tau\hat{H}_{\mathrm{SD}}}|n=0\rangle=Z_{R}(\tau,\bar{\mu},m_{R}).$ (72)
@@
$\langle\mathcal{O}(\tau)\mathcal{O}(0)\rangle_{\beta}$ $=\int_{0}^{\pi}d\bar{\alpha}\leavevmode\nobreak\ \rho(\bar{\alpha})\langle n_{L}=0|e^{-\tau\hat{H}_{m_{L},\bar{\mu}}}|n_{L}=0\rangle_{\mathrm{SC}}\,\langle n_{R}=0|e^{-\tau\hat{H}_{m_{R},\bar{\mu}}}|n_{R}=0\rangle_{\mathrm{SC}},$
$=\int_{0}^{\pi}d\bar{\alpha}\rho(\bar{\alpha})\int_{0}^{\pi}\int_{0}^{\pi}d\theta_{L}d\theta_{R}\leavevmode\nobreak\ \rho_{\Delta_{L},\bar{\alpha}}(\theta_{L})\rho_{\Delta_{R},\bar{\alpha}}(\theta_{R})\,e^{(\beta-\tau)\frac{\cos(\theta_{L})}{\hbar}}e^{\tau\frac{\cos(\theta_{R})}{\hbar}}$ (74)
$=\int_{0}^{\pi}\int_{0}^{\pi}d\theta_{L}d\theta_{R}\leavevmode\nobreak\ \rho(\theta_{L})\rho(\theta_{R})\,e^{(\beta-\tau)\frac{\cos(\theta_{L})}{\hbar}}e^{\tau\frac{\cos(\theta_{R})}{\hbar}}$
@@ 3.1 One-particle Hilbert space
$\langle\theta_{L}|e^{-\Delta\hat{L}}|\theta_{R}\rangle_{\mathrm{FC}}=\int_{0}^{\pi}d\bar{\alpha}\leavevmode\nobreak\ \rho(\bar{\alpha})\langle\theta_{L}|e^{-\Delta_{L}\hat{L}}|\bar{\alpha}\rangle_{\mathrm{FC}}\langle\bar{\alpha}|e^{-\Delta_{R}\hat{L}}|\theta_{R}\rangle_{\mathrm{FC}}.$ (77)
@@
For the identity operator in semi-open channel, as illustrated in Fig. 5(a), there is no time evolution along the boundary. Thus, a cut in the middle just prepares a state with $n=0$, reflecting the ”absence of structure” [19, 21]131313 As discussed in [19], setting $n_{M}=0$ implements a smooth gluing of the left and right boundaries. In the bulk, this can be understood as ensuring that the two split geodesics coincide, leaving no intermediate space between them.
@@
$|\Delta;\theta_{L},\theta_{R}\rangle=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=\int_{0}^{\pi}d\bar{\alpha}\rho(\bar{\alpha})\,|{\theta_{L}}\rangle_{\Delta_{L},\bar{\alpha}}\otimes^{\bar{\alpha}}|\bar{\alpha}\rangle\otimes^{\bar{\alpha}}|\theta_{R}\rangle_{\Delta_{R},\bar{\alpha}},$ (81)
@@
$\quad|\Delta;\theta_{L},\theta_{R}\rangle_{2}=\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}\,d\bar{\beta}\,\rho_{\Delta_{R_{1}},\bar{\beta}}(\bar{\alpha})\leavevmode\nobreak\ \rho(\bar{\beta})\leavevmode\nobreak\ |{\theta_{L}}\rangle_{\Delta_{L},\bar{\alpha}}\otimes^{\bar{\alpha}}|\bar{\alpha}\rangle_{\Delta_{R_{1}},\bar{\beta}}\otimes^{\bar{\beta}}|\bar{\beta}\rangle\otimes^{\bar{\beta}}|\theta_{R}\rangle_{\Delta_{R_{2}},\bar{\beta}}.$ (92)
Here, $\Delta=\Delta_{L}+\Delta_{R_{1}}+\Delta_{R_{2}}$. Note that we have also replaced $\rho(\bar{\alpha})$ with $\rho(\bar{\alpha})_{\Delta_{R_{1}},\bar{\beta}}$, since now the $\alpha$-region is twisted by EOW brane state $e^{-\Delta_{R_{1}}\hat{L}}|\bar{\beta}\rangle$ in the fully-open channel, as can be seen in (91). Accordingly, in the semi-open channel, one should expand $|n_{M_{1}}=0\rangle$ state in the twisted Hilbert space as: $|n_{M_{1}}=0\rangle=\int_{0}^{\pi}d\bar{\alpha}\,\rho(\bar{\alpha})_{\Delta_{R_{1}},\bar{\beta}}\leavevmode\nobreak\ |\bar{\alpha}\rangle_{\Delta_{R_{1}},\bar{\beta}}$.
@@
It may appear that the untwisted state $|\bar{\alpha}_{i}\rangle$ always occupies the second-to-last position in the decomposition. However, this is simply because at each iteration, we have inserted the new identity operator into the last $e^{-\Delta_{i}\hat{L}}$ in the fully-open channel. If we have arbitrary insertions, the untwisted state can be placed at any position in the tensor product. For example, the energy basis in the $n$-split representation can also be written as
@@
0-split $\begin{aligned} \\ \langle\theta_{L}|e^{-\Delta\hat{L}}|\theta_{R}\rangle\end{aligned}$ [Uncaptioned image]
$\begin{aligned} |{\theta_{L}}\rangle\otimes^{\theta_{L}}|{\theta_{R}}\rangle_{\Delta,\theta_{L}}\end{aligned}$ [Uncaptioned image]
1-split $\begin{aligned} \\ \int_{0}^{\pi}d\bar{\alpha}\rho(\bar{\alpha})\langle\theta_{L}|e^{-\Delta_{1}\hat{L}}|\bar{\alpha}\rangle\langle\bar{\alpha}|e^{-\Delta_{2}\hat{L}}|\theta_{R}\rangle\end{aligned}$ [Uncaptioned image]
$\begin{aligned} \int_{0}^{\pi}d\bar{\alpha}\,\rho(\bar{\alpha})\,|{\theta_{L}}\rangle_{\Delta_{1},\bar{\alpha}}\otimes^{\bar{\alpha}}|\bar{\alpha}\rangle\otimes^{\bar{\alpha}}|\theta_{R}\rangle_{\Delta_{2},\bar{\alpha}}\end{aligned}$
2-split $\begin{aligned} \\ &\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}_{1}d\bar{\alpha}_{2}\rho(\bar{\alpha}_{1})\rho(\bar{\alpha}_{2})\\ &\langle\theta_{L}|e^{-\Delta_{1}\hat{L}}|\bar{\alpha}_{1}\rangle\langle\bar{\alpha}_{1}|e^{-\Delta_{2}\hat{L}}|\bar{\alpha}_{2}\rangle\langle\bar{\alpha}_{2}|e^{-\Delta_{3}\hat{L}}|\theta_{R}\rangle\end{aligned}$ [Uncaptioned image] $\begin{aligned} &\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}_{1}d\bar{\alpha}_{2}\rho_{\Delta_{2},\bar{\alpha}_{2}}(\bar{\alpha}_{1})\rho(\bar{\alpha}_{2})\\ &|{\theta_{L}}\rangle_{\Delta_{1},\bar{\alpha}_{1}}\otimes^{\bar{\alpha}_{1}}|\bar{\alpha}_{1}\rangle_{\Delta_{2},\bar{\alpha}_{2}}\otimes^{\bar{\alpha}_{2}}|\bar{\alpha}_{2}\rangle\otimes^{\bar{\alpha}_{2}}|\theta_{R}\rangle_{\Delta_{3},\bar{\alpha}_{2}}\end{aligned}$
@@
$n$-split $\begin{aligned} &\prod_{i=1}^{n}\left(\int_{0}^{\pi}d\bar{\alpha}_{i}\rho(\bar{\alpha}_{i})\right)\\ &\langle\theta_{L}|e^{-\Delta_{1}\hat{L}}|\bar{\alpha}_{1}\rangle\langle\bar{\alpha}_{1}|\cdot\cdot\cdot|\bar{\alpha}_{n}\rangle\langle\bar{\alpha}_{n}|e^{-\Delta_{n}\hat{L}}|\theta_{R}\rangle\end{aligned}$ [Uncaptioned image] $\begin{aligned} \\ &\prod_{i=1}^{n-1}\left(\int_{0}^{\pi}d\bar{\alpha}_{i}\rho_{\Delta_{i+1},\bar{\alpha}_{i+1}}(\bar{\alpha}_{i})\right)\int_{0}^{\pi}d\bar{\alpha}_{n}\rho(\bar{\alpha}_{n})\\ &|{\theta_{L}}\rangle_{\Delta_{1},\bar{\alpha}_{1}}\otimes^{\bar{\alpha}_{1}}|\bar{\alpha}_{1}\rangle_{\Delta_{2},\bar{\alpha}_{2}}\otimes^{\bar{\alpha}_{2}}|\bar{\alpha}_{2}\rangle_{\Delta_{3},\bar{\alpha}_{3}}\otimes^{\bar{\alpha}_{3}}\cdot\cdot\cdot\\ &\otimes^{\bar{\alpha}_{n-1}}|\bar{\alpha}_{n-1}\rangle_{\Delta_{n},\bar{\alpha}_{n}}\otimes^{\bar{\alpha}_{n}}|\bar{\alpha}_{n}\rangle\otimes^{\bar{\alpha}_{n}}|\theta_{R}\rangle_{\Delta_{n+1},\bar{\alpha}_{n}}\end{aligned}$

### v2

@@ 2.2 EOW brane quantization
$K_{AdS}^{R}=m_{R},\quad\partial_{n_{\mathrm{AdS}}^{R}}\Phi_{\mathrm{AdS}}^{R}=\bar{\mu}+im_{R}e^{-i\Phi}.$ (45)
@@
$\mathds{1}=\int_{0}^{\pi}d\theta\penalty 10000\ \rho(\theta)|\theta\rangle\langle\theta|=\sum_{n=0}^{\infty}|n\rangle\langle n|,\quad\langle n|\theta\rangle=\frac{H_{n}(\cos(\theta)|q^{2})}{(q^{2};q^{2})_{n}^{1/2}}.$ (46)
The only unknown quantity is the wavefunction $\langle n|m_{R},\bar{\mu}\rangle$. As explained in [20], this wavefunction can be determined by expressing the effective mass of the brane in AdS space in two different ways. [^8] The first way is to directly match the parameter of EOW brane here to the JT gravity brane mass $m_{\mathrm{AdS}}$: $m_{\mathrm{AdS}}=\partial_{n_{\mathrm{AdS}}}\Phi_{\mathrm{AdS}}-\Phi_{\mathrm{AdS}}K_{\mathrm{AdS}}$. The second way is to use the Hamiltonian constraint to relate $m_{\mathrm{AdS}}$ to the JT momentum and extrinsic curvature. See [20] for details.
@@
where an operator ordering has been chosen. We can treat the above equation as a Schrodinger eigenvalue equation
@@
$H_{m_{R},\bar{\mu}}\rightarrow g^{-1}(L)H_{m_{R},\bar{\mu}}\penalty 10000\ g(L),\quad g(L)=\prod_{k=0}^{\infty}\left(1+2im_{R}e^{-(L+k\hbar)}-e^{-2(L+k\hbar)}\right)^{-1/2},$ (54)
@@
$2\cos(\theta){{}_{\Delta_{R},\bar{\alpha}}\langle}\theta|n\rangle=_{\Delta_{R},\bar{\alpha}}{\langle}\theta|n+1\rangle+(q^{2\Delta}e^{i\bar{\alpha}}+q^{2\Delta}e^{-i\bar{\alpha}})q^{n}{{}_{\Delta_{R},\bar{\alpha}}\langle}\theta|n\rangle+(1-q^{4\Delta}q^{n-1})(1-q^{n}){{}_{\Delta_{R},\bar{\alpha}}\langle}\theta|n-1\rangle.$ (60)
This is exactly the recursion relation for the Al Salam-Chihara polynomial $Q_{n}(\cos(\theta)|q^{2\Delta}e^{i\bar{\alpha}},q^{2\Delta}e^{-i\bar{\alpha}};q^{2})$. [^11] This wavefunction can also be directly recovered from the chord Hilbert space of the DSSYK by gauging the chord parity symmetry [41]. Taking into account the fact that the left and right eigenfunctions differ for a non-Hermitian Hamiltonian, the wavefunctions can be normalized as [20, 42]
@@
$\mathds{1}=\int_{0}^{\pi}d\theta\penalty 10000\ \rho_{\Delta_{R},\bar{\alpha}}(\theta)\,|\theta\rangle_{\Delta_{R},\bar{\alpha}}\,{{}_{\Delta_{R},\bar{\alpha}}}\langle\theta|=\sum_{n=0}^{\infty}|n\rangle\langle n|.$ (65)
@@ 2.3 Gluing condition and the two-point function
$=\int_{0}^{\pi}d\bar{\alpha}\penalty 10000\ \rho(\bar{\alpha})\penalty 10000\ \langle n=0|e^{-(\beta-\tau)\hat{H}_{\mathrm{SD}}}e^{-\Delta_{L}\hat{L}}|\bar{\alpha}\rangle_{\mathrm{FC}}\penalty 10000\ \langle\bar{\alpha}|e^{-\Delta_{R}\hat{L}}e^{-\tau\hat{H}_{\mathrm{SD}}}|n=0\rangle_{\mathrm{FC}},$ (70)
$=\int_{0}^{\pi}d\bar{\alpha}\penalty 10000\ \rho(\bar{\alpha})\penalty 10000\ \Psi_{L}(\beta-\tau,\bar{\alpha},\Delta_{L})\Psi_{R}(\tau,\bar{\alpha},\Delta_{R}),$
$=\int_{0}^{\pi}d\bar{\alpha}\penalty 10000\ \rho(\bar{\alpha})\penalty 10000\ Z_{L}^{\star}(\beta-\tau,\bar{\mu},m_{L})Z_{R}(\tau,\bar{\mu},m_{R}),$ (71)
@@
$\langle\bar{\alpha}|e^{-\Delta_{R}\hat{L}}e^{-\tau\hat{H}_{\mathrm{SD}}}|n=0\rangle=\langle\bar{\alpha}|q^{2\Delta_{R}\hat{n}}e^{-\tau\hat{H}_{\mathrm{SD}}}|n=0\rangle=\langle m_{R},\bar{\mu}|e^{-\tau\hat{H}_{\mathrm{SD}}}|n=0\rangle=Z_{R}(\tau,\bar{\mu},m_{R}).$ (72)
@@
$\langle\mathcal{O}(\tau)\mathcal{O}(0)\rangle_{\beta}$ $=\int_{0}^{\pi}d\bar{\alpha}\penalty 10000\ \rho(\bar{\alpha})\langle n_{L}=0|e^{-\tau\hat{H}_{m_{L},\bar{\mu}}}|n_{L}=0\rangle_{\mathrm{SC}}\,\langle n_{R}=0|e^{-\tau\hat{H}_{m_{R},\bar{\mu}}}|n_{R}=0\rangle_{\mathrm{SC}},$
$=\int_{0}^{\pi}d\bar{\alpha}\rho(\bar{\alpha})\int_{0}^{\pi}\int_{0}^{\pi}d\theta_{L}d\theta_{R}\penalty 10000\ \rho_{\Delta_{L},\bar{\alpha}}(\theta_{L})\rho_{\Delta_{R},\bar{\alpha}}(\theta_{R})\,e^{(\beta-\tau)\frac{\cos(\theta_{L})}{\hbar}}e^{\tau\frac{\cos(\theta_{R})}{\hbar}}$ (74)
$=\int_{0}^{\pi}\int_{0}^{\pi}d\theta_{L}d\theta_{R}\penalty 10000\ \rho(\theta_{L})\rho(\theta_{R})\,e^{(\beta-\tau)\frac{\cos(\theta_{L})}{\hbar}}e^{\tau\frac{\cos(\theta_{R})}{\hbar}}$
@@ 3.1 One-particle Hilbert space
$\langle\theta_{L}|e^{-\Delta\hat{L}}|\theta_{R}\rangle_{\mathrm{FC}}=\int_{0}^{\pi}d\bar{\alpha}\penalty 10000\ \rho(\bar{\alpha})\langle\theta_{L}|e^{-\Delta_{L}\hat{L}}|\bar{\alpha}\rangle_{\mathrm{FC}}\langle\bar{\alpha}|e^{-\Delta_{R}\hat{L}}|\theta_{R}\rangle_{\mathrm{FC}}.$ (77)
@@
For the identity operator in semi-open channel, as illustrated in Fig. 5(a), there is no time evolution along the boundary. Thus, a cut in the middle just prepares a state with $n=0$, reflecting the ”absence of structure” [19, 21] [^15] As discussed in [19], setting $n_{M}=0$ implements a smooth gluing of the left and right boundaries. In the bulk, this can be understood as ensuring that the two split geodesics coincide, leaving no intermediate space between them.
@@
$|\Delta;\theta_{L},\theta_{R}\rangle=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=\int_{0}^{\pi}d\bar{\alpha}\rho(\bar{\alpha})\,|{\theta_{L}}\rangle_{\Delta_{L},\bar{\alpha}}\otimes^{\bar{\alpha}}|\bar{\alpha}\rangle\otimes^{\bar{\alpha}}|\theta_{R}\rangle_{\Delta_{R},\bar{\alpha}},$ (81)
@@
$\quad|\Delta;\theta_{L},\theta_{R}\rangle_{2}=\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}\,d\bar{\beta}\,\rho_{\Delta_{R_{1}},\bar{\beta}}(\bar{\alpha})\penalty 10000\ \rho(\bar{\beta})\penalty 10000\ |{\theta_{L}}\rangle_{\Delta_{L},\bar{\alpha}}\otimes^{\bar{\alpha}}|\bar{\alpha}\rangle_{\Delta_{R_{1}},\bar{\beta}}\otimes^{\bar{\beta}}|\bar{\beta}\rangle\otimes^{\bar{\beta}}|\theta_{R}\rangle_{\Delta_{R_{2}},\bar{\beta}}.$ (92)
Here, $\Delta=\Delta_{L}+\Delta_{R_{1}}+\Delta_{R_{2}}$. Note that we have also replaced $\rho(\bar{\alpha})$ with $\rho(\bar{\alpha})_{\Delta_{R_{1}},\bar{\beta}}$, since now the $\alpha$-region is twisted by EOW brane state $e^{-\Delta_{R_{1}}\hat{L}}|\bar{\beta}\rangle$ in the fully-open channel, as can be seen in (91). Accordingly, in the semi-open channel, one should expand $|n_{M_{1}}=0\rangle$ state in the twisted Hilbert space as: $|n_{M_{1}}=0\rangle=\int_{0}^{\pi}d\bar{\alpha}\,\rho(\bar{\alpha})_{\Delta_{R_{1}},\bar{\beta}}\penalty 10000\ |\bar{\alpha}\rangle_{\Delta_{R_{1}},\bar{\beta}}$.
@@
It may appear that the untwisted state $|\bar{\alpha}_{i}\rangle$ always occupies the second-to-last position in the decomposition. However, this is simply because at each iteration, we have inserted the new identity operator into the last $e^{-\Delta_{i}\hat{L}}$ in the fully-open channel. If we have arbitrary insertions, the untwisted state can be placed at any position. For example, the energy basis in the $n$-split representation can also be written as
@@
0-split $\begin{aligned} \\ \langle\theta_{L}|e^{-\Delta\hat{L}}|\theta_{R}\rangle\end{aligned}$ [Uncaptioned image] $\begin{aligned} |{\theta_{L}}\rangle\otimes^{\theta_{L}}|{\theta_{R}}\rangle_{\Delta,\theta_{L}}\end{aligned}$ [Uncaptioned image]
1-split $\begin{aligned} \\ \int_{0}^{\pi}d\bar{\alpha}\rho(\bar{\alpha})\langle\theta_{L}|e^{-\Delta_{1}\hat{L}}|\bar{\alpha}\rangle\langle\bar{\alpha}|e^{-\Delta_{2}\hat{L}}|\theta_{R}\rangle\end{aligned}$ [Uncaptioned image] $\begin{aligned} \int_{0}^{\pi}d\bar{\alpha}\,\rho(\bar{\alpha})\,|{\theta_{L}}\rangle_{\Delta_{1},\bar{\alpha}}\otimes^{\bar{\alpha}}|\bar{\alpha}\rangle\otimes^{\bar{\alpha}}|\theta_{R}\rangle_{\Delta_{2},\bar{\alpha}}\end{aligned}$ [Uncaptioned image]
2-split $\begin{aligned} \\ &\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}_{1}d\bar{\alpha}_{2}\rho(\bar{\alpha}_{1})\rho(\bar{\alpha}_{2})\\ &\langle\theta_{L}|e^{-\Delta_{1}\hat{L}}|\bar{\alpha}_{1}\rangle\langle\bar{\alpha}_{1}|e^{-\Delta_{2}\hat{L}}|\bar{\alpha}_{2}\rangle\langle\bar{\alpha}_{2}|e^{-\Delta_{3}\hat{L}}|\theta_{R}\rangle\end{aligned}$ [Uncaptioned image] $\begin{aligned} &\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}_{1}d\bar{\alpha}_{2}\rho_{\Delta_{2},\bar{\alpha}_{2}}(\bar{\alpha}_{1})\rho(\bar{\alpha}_{2})\\ &|{\theta_{L}}\rangle_{\Delta_{1},\bar{\alpha}_{1}}\otimes^{\bar{\alpha}_{1}}|\bar{\alpha}_{1}\rangle_{\Delta_{2},\bar{\alpha}_{2}}\otimes^{\bar{\alpha}_{2}}|\bar{\alpha}_{2}\rangle\otimes^{\bar{\alpha}_{2}}|\theta_{R}\rangle_{\Delta_{3},\bar{\alpha}_{2}}\end{aligned}$ [Uncaptioned image]
@@
$n$-split $\begin{aligned} &\prod_{i=1}^{n}\left(\int_{0}^{\pi}d\bar{\alpha}_{i}\rho(\bar{\alpha}_{i})\right)\\ &\langle\theta_{L}|e^{-\Delta_{1}\hat{L}}|\bar{\alpha}_{1}\rangle\langle\bar{\alpha}_{1}|\cdot\cdot\cdot|\bar{\alpha}_{n}\rangle\langle\bar{\alpha}_{n}|e^{-\Delta_{n}\hat{L}}|\theta_{R}\rangle\end{aligned}$ [Uncaptioned image] $\begin{aligned} \\ &\prod_{i=1}^{n-1}\left(\int_{0}^{\pi}d\bar{\alpha}_{i}\rho_{\Delta_{i+1},\bar{\alpha}_{i+1}}(\bar{\alpha}_{i})\right)\int_{0}^{\pi}d\bar{\alpha}_{n}\rho(\bar{\alpha}_{n})\\ &|{\theta_{L}}\rangle_{\Delta_{1},\bar{\alpha}_{1}}\otimes^{\bar{\alpha}_{1}}|\bar{\alpha}_{1}\rangle_{\Delta_{2},\bar{\alpha}_{2}}\otimes^{\bar{\alpha}_{2}}|\bar{\alpha}_{2}\rangle_{\Delta_{3},\bar{\alpha}_{3}}\otimes^{\bar{\alpha}_{3}}\cdot\cdot\cdot\\ &\otimes^{\bar{\alpha}_{n-1}}|\bar{\alpha}_{n-1}\rangle_{\Delta_{n},\bar{\alpha}_{n}}\otimes^{\bar{\alpha}_{n}}|\bar{\alpha}_{n}\rangle\otimes^{\bar{\alpha}_{n}}|\theta_{R}\rangle_{\Delta_{n+1},\bar{\alpha}_{n}}\end{aligned}$ [Uncaptioned image]